Circular No. 5669

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMPTON GAMMA RAY OBSERVATORY SCHEDULE N. Gehrels, Goddard Space Flight Center, NASA, on behalf of the

Compton Gamma Ray Observatory science team, reports the following observing schedule as of Nov. 24 for Phase 2 of the mission (listed are starting dates and OSSE primary and secondary targets): 1992 Nov. 17, Her X-1, galactic plane (GP); Nov. 24, Her X-1, galactic center (GC); Dec. 1, 8, and 15, Cyg X-1, GC; Dec. 22 and 29, and 1993 Jan. 5, 3C 273, GP; Jan. 12, IC 4329A, PSR 1822–09; Feb. 2, NGC 4507, NGC 6814; Feb. 9, 2CG 010-31, 3C 390.3; Feb. 22, 1E 1740–2942, MCG –2-58-22; Feb. 25, 4U 0115+634, 47 Tuc; Mar. 9, WR 140, NGC 7314; Mar. 23, Crab pulsar, Mrk 509; Mar. 29, 1E 1740–2942, MCG –2-58-22; Apr. 1, Cen A, SN 1987A; Apr. 20, NGC 4151, MCG +8-11-11; May 4, 1E 1740–2942, MCG –2-58-22; May 7, SMC X-1, 3C 120; May 13, Crab pulsar, PSR 1951+32; May 25, NGC 4151, MCG +8-11-11; May 31, 1E 1740–2942, Cyg X-3; June 3, SMC X-1, 3C 120; June 15, GP, empty space; June 22, GP, Coma cluster; June 29 and July 8, GC, Coma cluster; July 13, Vela pulsar, Mrk 78; July 27, NGC 6814, NGC 4507; Aug. 3, QSO 1028+313, PSR 0740–28; Aug. 20, 1E 1740–2942, 3C 279. If planning coordinated observations, please contact E. Chipman (e-mail chipman@grossc.gsfc.nasa.gov or grossc::chipman; phone 301-286-7764; Code 668.1, NASA/GSFC, Greenbelt, MD 20771, USA).

(2060) CHIRON S. Larson and R. Marcialis, Lunar and Planetary Laboratory, obtained

Cousins red CCD images of (2060) Chiron with the Steward Observatory 2.3-m telescope on Nov. 30.4 UT that showed a well-developed tail at least 50′′ long (possibly off the field at 80′′) at p.a. 288^{o}.

NOVA IN THE LARGE MAGELLANIC CLOUD 1992 Additional photometry by A. C. Gilmore at Mt. John (cf. IAUC 5661):

Nov. 27.55 UT, V = 13.06 ± 0.02, U–B = –0.72 ± 0.04, B–V = +0.02 ± 0.02, V–R = +0.92 ± 0.01 (reference star CPD −65^{o}433).

PERIODIC COMET KOJIMA (1992z)

Corrigendum. On IAUC 5667, line 4, for = –0.05 day read < |0.01| day

1992 December 5 Daniel W. E. Green